### v1

# Novel Topological Insulators with Hybrid-order Boundary States

Abstract
We report the discovery of several classes of novel topological insulators (TIs) with hybrid-order boundary states generated from the first-order TIs with additional crystalline symmetries. Unlike the current studies on hybrid-order TIs where different-order topology arises from merging different-order TIs in various energy, these novel TIs exhibit a remarkable coexsitence of first-order gapless modes and higher-order Fermi-arc states, behaving as a hybrid between the first-order TIs and higher-order topological semimetals within a single bulk gap. Our findings establish a profound connection between these novel $d$-dimensional ($d$D) TIs and ($d-1$)D higher-order TIs (HOTIs), which can be understood as a result of stacking $(d-1)$D HOTIs to $d$D with $d=3,4$, revealing unconventional topological phase transitions by closing the gap in certain first-order boundaries rather than the bulk. The bulk-boundary correspondence between these higher-order Fermi-arcs and bulk topological invariants acossiated with additional crystallline symmetries is also demonstrated. We then discuss the conventional topological phase transitions from these novel TIs to nodal-line/nodal-surface semimetal phases, where the gapless phases host new kinds of topological responses. Meawhile, we present the corresponding topological semimetal phases by stacking these unique TIs. Finally, we discuss potential ways to realize these novel phases in synthetic and real materials, with a particular focus on the feasible implementation in optical lattices using ultracold atoms.
Introduction.— Topological insulators (TIs) and topological semimetals (TSMs) have emerged as one of the most active fields of modern physics over the past two decades Hasan and Kane (2010); Qi and Zhang (2011); Armitage et al. (2018), attracting significant attention from researchers in condensed matter physics and artificial systemsZhang et al. (2018); Cooper et al. (2019); Xu (2019); Ozawa et al. (2019); Lu et al. (2014); Liu et al. (2020); Zhu et al. (2023). One significant property of topological phases is the presence of bulk-boundary correspondence, which guarantees the existence of gapless first-order boundary modes that are associated with certain bulk topological invariant. These first-order topological materials have been classifed within the framework provided by the real K-theory Atiyah (1966) and Altland-Zirnbauer (AZ) classes Chiu et al. (2016) for both gapped Schnyder et al. (2008); Kitaev (2009) and gapless Hořava (2005); Zhao and Wang (2013, 2014) systems, based on their fundamental symmetries including time-reversal $\mathcal{T}$, charge-conjugation $\mathcal{C}$, and chiral symmetry $\mathcal{S}$.
Recent research focus has been extended to higher-order topological phases with additional crystalline symmetries Benalcazar et al. (2017); Song et al. (2017); Langbehn et al. (2017); Schindler et al. (2018a, b). These systems do not feature first-order even second-order boundary gapless states but instead exhibit gapless states at higher-order boundaries. For instance, a second-order TI (SOTI) in 3D hosts boundary gapless modes at 1D hinges (i.e., second-order boundary) while its 2D surface (i.e., first-order boundary) spectrum is gapped. By further gapping the gapless hinge states, a 3D third-order TI (TOTI) can be induced, leading to the existence of zero-energy modes at 0D corners (i.e., third-order boundary). By stacking these HOTIs along an extra dimension, one can obtain higher-order TSMs (HOTSMs) with higher-order Fermi-arcs Wang et al. (2020a); Ghorashi et al. (2020); Simon et al. (2022); Wei et al. (2021). Moreover, hybrid-order TIs Kooi et al. (2020) with multi-gaps have been also discovered in metamaterials Zhang et al. (2020); Yang et al. (2021); Lai et al. (2023), which simultaneously host hybrid-order topological boundary modes under different open boundary conditions (OBCs) in various energy. Building on recent advancements in the aforementioned areas, a natural question arises: Are there novel types of TIs that host hybrid-order boundary modes within a single bulk gap? Can we establish bulk-boundary correspondence in these phases?
In this Letter, we answer these questions positively. We introduce several novel classes of TIs that exhibit unique hybrid-order boundary states consisting of both first-order gapless modes and higher-order Fermi-arc states within a single bulk gap, which can be seen as a hybrid between first-order TIs (FOTIs) and HOTSMs. These phases can be simply generated from the first-order TIs (FOTIs) by shifting the gapless modes in first-order boundary Brillouin Zone (BZ) through specific perturbations protected by crystalline symmetries. On the other hand, we can also regard these unique TI phases in $d$D as the stacking $(d-1)$D HOTIs along an additional spatial dimension ($d=3,4$). In this viewpoint, we reveal unconventional topological phase transitions (UTPTs) by closing the first-order boundary gap while maintaining an open bulk gap, i.e., the unconventional phase transition points are topological crystalline insulators (TCIs) Fu (2011). We emphasize these higher-order Fermi-arcs can be well described by the bulk topology associated with related crystalline symmetries. For concreteness, we mainly focus on several 4D models in the main text: (i) Class A with combined mirror symmetry; (ii) Class AIII with $\mathcal{RT}$-symmetry ; (iii) Class A with $C_{4}\mathcal{T}$-symmetry; (iv) Class AIII with $C_{4}\mathcal{T}$-symmetry. In (i), we find a 4D TI hosts both first-order boundary Weyl cones and second-order chiral hinge Fermi-arcs located at the typical hinges determined by the related mirror reflection symmetries; In (ii), a 4D TI with spin conservation hosts first-order boundary Dirac cones and second-order helcial hinge Fermi-arcs while Dirac cones will be expanded into real Dirac nodal-lines and helical hinge states will be gapped leading to third-order Fermi-arcs after introducing a spin-orbit-coupling (SOC) $\mathcal{RT}$-symmetry perturbation; In (iii), a 4D TI with $C_{4}\mathcal{T}$ is similar to the case (i) but with the difference that the second-order hinge Fermi-arc are always located at four hinges determined by more combined mirror-symmetries; (iv) Two copies of Hamiltonian with opposite sign in (iii) are considered which supports first-order boundary Dirac cones and second-order helical hinge Fermi-arcs. By performing dimensional reduction with $k_{z}=0$, the corresponding 3D models and physical pictures are obtained, which are presented in the Supplemental Materials (SM). The explicit expressions of the topological numbers associated with the corresponding crystalline symmetries are presented. Moreover, we also explore the conventional topological phase transition (CTPT) from a novel TI in case (i)/(ii) to a nodal-line/nodal-sphere phase, presenting the topological response theory of these gapless phases. Furthermore, we discuss the related topological semimetal phases by stacking these novel TIs along an additional spatial dimension. Finally, we discuss the possible ways to realize these novel phases in synthetic and real materials, including the implemented proposals in optical lattices using ultracold atoms.
[…]
$\begin{split}&\mathcal{H}_{1}(\bm{k})=\mathcal{H}_{0}+\Delta_{1},\\ \end{split}$ (1)
where $\mathcal{H}_{0}=\sum_{i}d_{i}\Gamma_{i}+d_{0}\Gamma_{0}$ with the Bloch vector $d_{i}=\sin k_{i}$, $d_{0}=M-\sum_{i}\cos k_{i}$ ($i=x,y,z,w$). The perturbation $\Delta_{1}=b_{xw}\Gamma_{0}\Gamma_{14}+b_{yw}\Gamma_{0}\Gamma_{24}$. Here the Driac matrices are $\Gamma_{1}=G_{31}$, $\Gamma_{2}=G_{32}$, $\Gamma_{3}=G_{33}$, $\Gamma_{4}=G_{20}$, $\Gamma_{0}=G_{10}$, $\Gamma_{ab}=\frac{i}{2}[\Gamma_{a},\Gamma_{b}]$, and satisfy $\{\Gamma_{i},\Gamma_{j}\}=2\delta_{ij}$. We mix the notation $i\in(1,2,3,4)\leftrightarrow(x,y,z,w)$ in the subscript of $\Gamma_{i}$ and label $G_{ab}=\sigma_{a}\otimes\sigma_{b}$ through the $2\times 2$ identity $\sigma_{0}$ and Pauli matrices $\sigma_{i}$ hereafter. As one can see that $\mathcal{H}_{0}$ is the well-known 4D quantum Hall insulator (QHI) characterized by the second Chern number (SCN) $C_{2}$ SCN , harboring $|C_{2}|$ Weyl cones at the origin of the 3D boundary BZ Qi et al. (2008). Under OBCs along the $x$ ($w$)-direction with a chain length $L$, introducing a non-zero $b_{xw}$ term in $\mathcal{H}_{0}$ leads to the shifted Weyl cones along the $k_{w}$ ($k_{x}$)-axis in the boundary BZ, resulting in the second-order chiral Fermi-arc hinge states when considering OBCs at $x$ and $y$ directions. A similar effect is observed when only the $b_{yw}$ term is non-zero, as shown in Fig. 1. We present the effective boundary Hamiltonian and numeric results in the SM SM .
In what follows we discuss the topology of this system. The initial Hamiltonian $\mathcal{H}_{0}(\bm{k})$ in class AII respects $\mathcal{T}$ represented as $\hat{\mathcal{T}}=\Gamma_{13}\mathcal{K}$, where $\mathcal{K}$ is the complex conjugate. In addition, $\mathcal{H}_{0}$ preserves extra crystalline symmetries, including inversion symmetry $\mathcal{I}$ IS and the combined mirror reflection symmetries $\mathcal{M}_{ij}=\mathcal{M}_{i}\mathcal{M}_{j}$, where each mirror reflection $\mathcal{M}_{i}$ inverts one of the momentum components $k_{i}$ MS . We present all the definitions of the corresponding crytalline symmetries in the SM SM . For simplicity, we first consider the case when only $b_{xw}$ term in $\Delta_{1}$ is numerically small, it breaks $\mathcal{T}$ and only preserves $\mathcal{M}_{xw}$ and $\mathcal{M}_{yz}$ symmetries. The system now falls into class A and still hosts a non-trivial $C_{2}$ with the shifted first-order Weyl boundary states. To characterize the second-order topology, we can define extra topological numbers at $\mathcal{M}_{xw}$/$\mathcal{M}_{yz}$-invariant points $\Lambda_{a}^{xw}/\Lambda_{a}^{yz}$ HSP , i.e., named “mirror first Chern number” (MFCN),
$C_{1m}^{ij}(\Lambda_{a}^{ij})=[C_{+i}(\Lambda_{a}^{ij})-C_{-i}(\Lambda_{a}^{ij})]/2,$ (2)
where $C_{\pm i}$ denotes the first Chern number defined in 2D subsystems $\mathcal{H}_{1}^{\Lambda_{a}^{ij}}(k_{l},k_{m})$ of the eigenspace =±$\mathcal{E}=\pm i$ of ^$\hat{\mathcal{M}}_{ij}$ MS . For simplicity and without loss of generality, we focus on the case when $M$ near $3$ hereafter, evaluation shows that $C_{1m}^{xw}=-1$/$C_{1m}^{yz}=1$ at $\Lambda_{1}^{xw}/\Lambda_{1}^{yz}=(0,0)$, and $C_{1m}^{xw}(\Lambda_{a}^{xw})=0$/$C_{1m}^{yz}(\Lambda_{a}^{yz})=0$ for the rest of the three $\Lambda_{a}^{xw/yz}$ points SM . These results establish the bulk-hinge correspondence and provide detailed information of the chiral hinge Fermi-arcs, as depicted in Fig. 1(a). Since these hinge Fermi-arcs are formed by stacking the 1D chiral modes $\mathcal{H}_{R/L}=\pm k_{z}$ along $k_{w}$ axis, once $\mathcal{M}_{xw}$ acts on one of hinge states, it will transform this states to the mirror reflection hinge about $x=0$ and $k_{w}=0$ planes without changing its chirality; while $\mathcal{M}_{yz}$ tranforms one of the hinge states to its mirror-symmetric hinge about $y=0$ and then changes the chirality of this hinge mode due to the mirror reflection about $k_{z}=0$ plane, causing ±$\pm k_{z}\rightarrow\mp k_{z}$. Similar picture also appear when $M$ is in other parameter regionsSM . When only $b_{yw}$ is non-zero, the system preserves both $\mathcal{M}_{xz}$ and $\mathcal{M}_{yw}$ symmetries describing the existence of second-order chiral Fermi-arcs, as depicted in Fig. 1 (b).
When both $b_{xw}$ and $b_{yw}$ survive, $\mathcal{M}_{xw}$ and $\mathcal{M}_{yz}$ are broken. For the typical case when $b_{xw}=b_{yw}$, $\mathcal{H}_{1}$ hosts an extra combined symmetry $\mathcal{M}_{(x-y)z}=\mathcal{M}_{x-y}\mathcal{M}_{z}$ with $\mathcal{M}_{x-y}$ being the mirror reflection about the mirror-invariant line $k_{x}=k_{y}$ (3). In the same spirit, we can calculate the MFCNs $C_{1m}^{k_{z}}$ at $k_{x}=k_{y}$, $k_{z}=0,\pi$ planes with $C_{1m}^{0}=-1$ and $C_{1m}^{\pi}=0$SM . As we can see that chiral Fermi-arcs are located at diagonal hinges with mirror reflection about $x=y$ plane harboring opposite chirality due to the operation of $\mathcal{M}_{z}$, as shown in Fig. 1(c). Similarly, when $b_{xw}=-b_{yw}$, $\mathcal{H}_{1}$ preserves $\mathcal{M}_{(x+y)z}$ symmetry with $C_{1m}^{0}=1$ and $C_{1m}^{\pi}=0$, enforcing the location of Fermi-arcs at anti-diagonal corners with opposite chiralitySM . If $b_{xw}\neq|b_{yw}|$, $\Delta_{1}$ breaks $\mathcal{T}$ and all the $\mathcal{M}_{ij}$-symmetries. However, the system still hosts the $\mathcal{I}$-symmetry which guarantees the second-order Fermi-arc states as a superposition of the cases where either $b_{yw}=0$ or $b_{xw}=0$. We emphasize that $\mathcal{M}_{ij}$-symmetry is used to construct this unique TI; it is itself not essential for the existence of second-order hinge Fermi-arcs. The hinge states are robust against a $\mathcal{M}_{ij}$-symmetry breaking perturbation, as long as the bulk gap is not closed.
Figure 1: Schematic of the boundary states of $\mathcal{H}_{1}$ with $C_{2}=-1$ when $42<M<$ 4. Each first-order 3D boundary hosts a 3D Weyl cone marked in red/blue carrying right-/left- chirality while the 1D chiral hinge modes are also marked in the same way. (a) $b_{xw}$ term shifts the Weyl cone along the $k_{w}$ direction, causing it to be located at $\mathbf{k}_{x^{+}}=(k_{y},k_{z},k_{w})=(0,0,b_{xw})$ at $x=-L/2$, while the Weyl cone at $x=L/2$ is shifted and located at $\mathbf{k}_{x^{-}}=(0,0,-b_{xw})$. If we further open the boundary along the $y$ axis, there will be 1D chiral zero-energy Fermi-arcs located at the symmetry-protected hinges for $k_{z}=0$. (b) $b_{yw}$ term shifts the Weyl cone at $y=-L/2$ to the position $\mathbf{k}_{y^{+}}=(k_{x},k_{z},k_{w})=(0,0,b_{yw})$ and another Weyl cone at $y=L/2$ to $\mathbf{k}_{y^{-}}=(0,0,-b_{yw})$. 1D chiral hinge Fermi-arcs also appear at chosen hinges when OBC along $x$ and $y$ directions are assumed. (c) The superposition of the cases in Figs. (a) and (b) with $b_{xw}=b_{yz}$ where second-order Fermi-arcs only appear at two $\mathcal{M}_{(x-y)z}$-symmetric hinges. (d)Phase diagram of 3D subsystems $\mathcal{H}^{k_{w}}_{1}$ with only $b_{xw}$ being non-zero by varying $k_{w}$. Here“NI” denotes the normal insulator.
To understand this 4D novel phase with both first- and second-order boundary states, we can treat $k_{w}$ as a parameter in this $\mathcal{H}_{1}$. The 3D subsystem $\mathcal{H}_{1}^{k_{w}}(k_{x},k_{y},k_{z})$ goes through UTPTs while the gap closing happens in certain 2D surface BZs instead of the 3D bulk when tuning $k_{w}$ from $-\pi$ to $\pi$. For instance and when $b_{xw}>0$, the 3D subsystems exhibit two distinct phases, which are normal insulators for $k_{w}\in(-\pi,-b_{xw})\cup(b_{xw},\pi)$, and chiral SOTIs for $k_{w}\in(-b_{xw},0)\cup(0,b_{xw})$. The TCIs Fu (2011) at $k_{w}=0,\pm b_{xw}$ serve as the unconventional transition points between these two phases, as illustrated in Fig. 1(d). Thus this novel 4D TI phase $\mathcal{H}_{1}$ is nothing but the stacking layers of 3D chiral SOTIs, revealing UTPTs by tuning $k_{w}$. We present detailed discussions about the topology of these 3D subsystems in the SMSM . On the other hand, a 3D novel TI with hybrid-order boundary states can be simply obtained by setting $k_{z}=0$, the system restores $\mathcal{S}$-symmetry and thus falls into class AIII. Similar discussions can be found in the SMSM .
Eight-band model in Class AIIIRT without and with SOC.— We now consider a 4D eight-band model, its Hamiltonian is given by,
$\begin{split}\mathcal{H}_{2}(\bm{k})=\tilde{\mathcal{H}}_{0}+\Delta_{2}+\Delta_{3},\end{split}$ (3)
where $\tilde{\mathcal{H}_{0}}=\sum_{i}d_{i}\tilde{\Gamma}_{i}+d_{0}\tilde{\Gamma}_{0}$ with the Bloch vector $d_{i}$ $(i=x,y,z,w)$ are the same as in $\mathcal{H}_{0}$, the perturbations $\Delta_{2}=b_{xw}\tilde{\Gamma}_{0}\tilde{\Gamma}_{14}+b_{yw}\tilde{\Gamma}_{0}\tilde{\Gamma}_{24}+b_{zw}\tilde{\Gamma}_{0}\tilde{\Gamma}_{34}$, and $\Delta_{3}=c_{x}\tilde{\Gamma}_{61}+c_{z}\tilde{\Gamma}_{63}$. Here $\tilde{\Gamma}_{1}=G_{331}$, $\tilde{\Gamma}_{2}=G_{332}$, $\tilde{\Gamma}_{3}=G_{333}$, $\tilde{\Gamma}_{4}=G_{320}$, $\tilde{\Gamma}_{0}=G_{310}$, $\tilde{\Gamma}_{6}=G_{100}$, $\tilde{\Gamma}_{7}=G_{200}$, and $\tilde{\Gamma}_{ab}=\frac{i}{2}[\tilde{\Gamma}_{a},\tilde{\Gamma}_{b}]$. Label $G_{ijk}=\sigma_{i}\otimes\sigma_{j}\otimes\sigma_{k}$ hereafter. ~$\tilde{\mathcal{H}}_{0}$ is just a 4D quantum spin Hall insulator as two copies of 4D QHIs ±$\pm\mathcal{H}_{0}$ with opposite ±$\pm C_{2}$ implying the existence of 3D first-order boundary Dirac modes. $\Delta_{2}$ shifts the gapless Dirac cones in 3D boundary momentum space similar to that case in $\mathcal{H}_{0}$, resulting in second-order helical Fermi-arc hinge states. $\Delta_{3}$ behaves like a SOC, where $c_{i}$ expands the 3D Dirac cone into a nodal line structure at $k_{i}=0$ plane at the boundary normal to $j$-direction. For instance, non-zero $c_{x}$ term expands the boundary Dirac cone into a nodal line at $k_{x}=0$ plane in the case of OBC at $y/z/w$-directionSM .
In the following we discuss the topology of this model in details. In AZ classification Chiu et al. (2016), ~$\tilde{\mathcal{H}}_{0}$ falls into class CII with $\mathcal{T}$, $\mathcal{C}$ and $\mathcal{S}$ hosting a $\mathbb{Z}_{2}$ topologyN (4). Since ~$\tilde{\mathcal{H}}_{0}$ is spin conservation with $\tilde{\Gamma}_{5}$-symmetry, we have $[\tilde{\Gamma}_{5},\tilde{\mathcal{H}}_{0}]=0$, where $\hat{\tilde{\Gamma}}_{5}=G_{300}$. We can define a $\mathbb{Z}_{2}$ number through the spin SCN N (5) to characterize the first-order topology with bulk-boundary correspondence even if $\mathcal{T}$ and $\mathcal{C}$ will be broken laterZhu et al. (2022). In addition, ~$\tilde{\mathcal{H}}_{0}$ also respects the following crystalline symmetries, including $\mathcal{I}$, and $\mathcal{M}_{i}$ leading to the combined symmetries $\mathcal{M}_{ij}$ and $\mathcal{R}_{i}\mathcal{T}=\mathcal{M}_{i}\mathcal{I}\mathcal{T}$ for $i=x,y,z,w$N (6).
Figure 2: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4. (a-b)$b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}=0$, the system harbors both shifted first-order Dirac points located at the same positions as in Fig. 1(c) and second-order Fermi-arcs consisting of helical hinge states (colored in orange) of 3D SOTIs stacked along $k_{w}$ axis. (c-e) $b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}>0$. The system harbors both shifted first-order nodal lines and third-order Fermi-arcs consisting of zero-mode corner states of 3D TOTIs stacked along $k_{w}$ axis.
We first consider the situation when $\Delta_{2}$ plays the role in, i.e., all $c_{i}=0$. $\Delta_{2}$ break $\mathcal{T}$, $\mathcal{C}$, but keep $\mathcal{S}$, $\mathcal{I}$ and thus the system is now falls into calss AIII but still hosts an unchanged spin SCN. The discussion of the second-order topology associated with the spin MFCNs N (8) is similar to the cases in $\mathcal{H}_{1}$ when only $b_{xw}$ and $b_{yw}$ surviveSM . To conveniently discuss third-order topology later, we focus on a TI phase with both first- and second-order boundary states with $b_{xw}=b_{yw}=b_{zw}$ for $M$ near $3$. As shown in Fig. 2(a-b), the system hosts the second-order helical Fermi-arcs where its 3D layers $\mathcal{H}_{2}^{k_{w}}$ with $k_{w}\in(-b_{xw},b_{xw})$ are SOTIs with helical states propagating along $\mathcal{M}_{x-y}$, $\mathcal{M}_{y-z}$, and $\mathcal{M}_{z-x}$ symmetric hingesN (1) when considering OBCs along $x$-, $y$- and $z$-directions. One can also define 3D mirror winding numbers to describe this second-order topology accordinglySM .
Subsequently, we turn on $c_{x}=c_{z}$ being numeric small, $\mathcal{H}_{3}$ remains in class AIII harboring shifted 3D $\mathbb{Z}_{2}$ nodal lines Shao and Zhao (2018); Wang et al. (2020b); Ahn et al. (2018) on the 3D boundaries when assuming OBC along $x$ or $y$ directions [Fig. 2(c)]. However, helical hinge Fermi-arcs have been gapped due to $\Delta_{3}$ breaks $\tilde{\Gamma}_{5}$ ($\mathcal{M}_{x-y}$ and $\mathcal{M}_{y-z}$ are also broken), leading to third-order Fermi-arcs along $k_{w}$ axisSM , as shown in Fig. 2(c-e). In this case, even though the spin SCN is not well-defined, one can define a $\mathbb{Z}_{2}$ number as 4D generalization of Kane-Mele invariant Shao and Zhao (2018) since $\mathcal{H}_{3}$ remains respect $\mathcal{R}_{w}\mathcal{T}$-symmetryN (6). Alternately, we can define another topological number as $\nu=\nu_{0}-\nu_{\pi}~{}\text{mod}~{}2$ at $\mathcal{R}_{w}\mathcal{T}$-invariant lines $k_{w}=0,\pi$ with $\nu_{k_{w}}$ is a topological number of the 3D real TI with $\mathcal{S}$ Dai et al. (2021). Calculation results show that $\nu_{0}=1$ and $\nu_{\pi}=0$ for $M$ near $3$ SM . Meanwhile, $\mathcal{H}_{3}$ keeps $\mathcal{M}_{(z+x)yw}=\mathcal{M}_{z+x}\mathcal{M}_{y}\mathcal{M}_{w}$ (7), $\mathcal{M}_{z-x}$, and $\mathcal{I}$-symmetries. A non-trivial 1D mirror winding number $w_{1m}=2$ defined at $k_{x}=-k_{z}$ and $(k_{y},k_{w})=(0,0)$ associated with $\mathcal{M}_{(z+x)yw}$-symmetry SM guarantees the exact location of third-order Fermi-arc with $\mathcal{M}_{w}$-symmetry about $k_{w}=0$ consisting of a pair of zero-modes, accumulating at two $\mathcal{M}_{(z+x)y}$-symmetric [combined mirror reflection about $z=-x$ and $y=0$ planes] corners in a 3D real space cubic spanned in $(x,y,z)$. The bulk-corner correspondence is established here; $\mathcal{M}_{z-x}$ implies these two zero-modes live in $z=x$ plane; these Fermi-arc states are always $\mathcal{I}$-symmetric in $(x,y,z,k_{w})$ space; see Fig. 2(d-e). We emphasize that $\tilde{\Gamma}_{5}$ $(\mathcal{R}_{w}\mathcal{T})$ symmetry guarantees the existence of second (third)-order gapless Fermi arcs in class AIII, where these boundary modes are located at the positions determined by additional crystalline symmetries.
Figure 3: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4 and $\delta>0$ being numeric small. (a) The system harbors both shifted first-order Weyl cones SM and the second-order chiral Fermi-arc hinge states. The green dashed squares denote the 3D TCIs being the unconventional phase transition points. (b) 3D non-trivial subsystems with fixing $k_{w}$ host second-order chiral hinge states.
[…]
$\mathcal{H}_{3}(\bm{k})=\mathcal{H}_{0}+\Delta_{4},$ (4)
with a perturbation $\Delta_{4}=\delta(\cos k_{x}-\cos k_{y})\Gamma_{4}$ breaks $C_{4}^{zw}$ and $\mathcal{T}$ individually, but respects $C_{4}^{zw}\mathcal{T}$-symmetry CT . As shown in Fig. 3, $\Delta_{4}$ shifts the boundary Weyl cones along $k_{w}$-direction SM . Two Weyl cones located on different boundaries [e.g. $y(x)=-L/2$ and $y(x)=L/2$] are shifted along the same direction when OBC along $y$ ($x$)-direction is assumed, resulting in a second-order Fermi-arc at four hinges once we further open the boundary along $x$ $(y)$ axis. Except the unchanged $C_{2}$ for the first-order topolgy, we can define the so-called “rotational spin first Chern number” (RSFCN) at $C_{4}^{zw}\mathcal{T}$-invariant 2D momenta $(k_{x},k_{y})=(0,0),(\pi,\pi)$ which may be used to characterize the second-order topology SM . In addition, we can easily construct the spin conservation verion of $\mathcal{H}_{3}$ with eight-band as $\mathcal{H}_{4}(\bm{k})=\sigma_{3}\otimes\mathcal{H}_{3}$, which respects $C^{zw}_{4}\mathcal{T}$ and $\mathcal{S}$ and thus falls into class AIII. The system now hosts the first-order boundary Dirac modes and the second-order helical hinge Fermi-arcs. By setting $k_{z}=0$, we obtain the corresponding 3D unique TI. All these similar discussions about the second-order topology associated with additional crystalline symmetries are presented in Ref.SM .
CTPTs and Novel TSMs.—By enlarging the parameters in $\Delta_{1}/\Delta_{2}$ of $\mathcal{H}_{1}$/$\mathcal{H}_{2}$, we can study the CTPT goes through a phase transition from this unique TI phase to a non-trivial nodal-line/sphere metallic phase and then finally becomes a trivial insulator. Such 4D nodal structures, including nodal-line, spin-nodal-line, and nodal-sphere, are characterized by the first Chern number (FCN) $C_{1}$, spin FCN $C_{1s}$, and 1D winding number $w_{1}$ respectively. The corresponding topological responses based on current works Wang et al. (2021); Ramamurthy and Hughes (2017); Hirsbrunner et al. (2023) are also discussed SM . On the other hand, we can obtain serveral new classes of TSMs with hybrid-order boundary Fermi-arcs by stacking the above TI phases along an extra spatial dimension $u$, as is commonly done, i.e., $M\rightarrow M^{\prime}=M-\cos k_{u}$, For instance, when $M=4$, $\mathcal{H}_{1}(M^{\prime})$/$\mathcal{H}_{1}^{k_{z}=0}(M^{\prime})$ hosts a pair of 5D/4D nodal surfaces/lines characterized by double charges $\nu_{D}=(C_{2}/w_{3},C_{1})$, where $w_{3}$ denotes the 3D winding number; $\mathcal{H}_{3}(M^{\prime})$/$\mathcal{H}_{3}^{k_{z}=0}(M^{\prime})$ hosts a pair of hybrid-order 5D/4D monopoles characterized by $C_{2}/w_{3}$. Two spin copies of these nodal-structures of $\mathcal{H}_{2}(M^{\prime})/\mathcal{H}_{4}(M^{\prime})$ can be also extended accordinglySM .
Discussion and outlook.— In this work, we establish a valuable framework for understanding the origin of the distinctive properties of these novel TIs, demonstrating that these unique TIs can be seen as a higher-dimensional stacking effect of low-dimensional HOTIs undergoing UTPTs along an additional spatial dimension. In other words, each kind of $d$D HOTI corresponds to a $(d+1)$D unique TI phase for $d\geq 2$. The related nodal structures and their topological responses are also studied when $\mathcal{H}_{1}/\mathcal{H}_{2}$ becomes a gapless phase. Furthermore, the corresponding TSMs with novel nodal-structures can be obtained by stacking these TIs going through CTPTs along an extra spatial dimension. Note that these unique topological phases, especially for models with $d\geq 4$, could be experimental studied in sythetic matter, as could be realized in metamaterialsCheng et al. (2021); Ma et al. (2021); Chen et al. (2021a, b), in electric circuitsWang et al. (2020c, 2023), or in optical lattices with cold atomsPrice et al. (2015); Petrides et al. (2018); Lohse et al. (2018); Bouhiron et al. (2024), etc. In particular, we can utilize the cold-atom setup proposed in our previous work Zhu et al. (2022) to realize the 4D model ~$\tilde{\mathcal{H}}_{0}$ in Eq. (3), including the implementation of other models mentioned in this work in a similar manner SM . For the 3D model in $\mathcal{H}_{1}$ with $k_{z}=0$, $\mathcal{H}_{0}^{k_{z}=0}$ as a 3D strong TI in low-energy limit has been predicted Zhang et al. (2009) and realized Chen et al. (2009) in real materials, such as Bi2Se3 and Bi2Te3 families, one can consider to induce $\Delta_{1}$ term in these materials by magnetization or light-induced Floquet drivings N (10).
We highlight that these novel topological phases can also be generalized to non-Hermitian Gong et al. (2018); Yao and Wang (2018) and Floquet systems Cayssol et al. (2013); Goldman and Dalibard (2014). The field-theoretical descriptions of these phases in the bulk is one of interesting directions. An intriguing and open question is the direct calculation of the effective action in the bulk. Take $\mathcal{H}_{1}$ as an example, the presence of shifted first-order boundary Weyl cones implies the possibility of an additional term beyond the $(4+1)$D Chern-Simons action in its 4D parent system. We provide a relevant discussion in the SM SM and will address it in the future. Overall, our work further open the door of investigating novel topological phases in the synthetic and real materials.
Acknowledgements.
This work was supported by the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the NSFC/RGC JRS grant (Grant No. N  HKU774/21), the CRF (Grant No. C6009-20G) and GRF (Grants No. 17310622 and No. 17303023) of Hong Kong, Guangdong Provincial Quantum Science Strategic Initiative (GDZX2200001), and National Key R&D Program of China (Grant No. 2022YFA1405304).